OpenStudy (anonymous):

lim x -> 0 of x csc x +1 / x csc x

OpenStudy (anonymous):

\[xcsc(x) + \frac{ 1 }{ xcsc(x) } = \frac{ 1 }{ \sin(x)/x }+ \frac{ sinx }{ x }\] Would you happen to know what lim x ->0 sinx/x is? :)

OpenStudy (paxpolaris):

\[\lim_{x \to 0} {\sin x \over x}=1\]